OpenStudy (anonymous):

Anti-derivative of (x^2+3x-4)/x ?

OpenStudy (anonymous):

split it up

OpenStudy (anonymous):

bobobo

OpenStudy (anonymous):

\[\int\left(\frac{x^2+3x-4}{x}\right)dx=\int\left(x+3-\frac{4}{x}\right)dx=\cdots\]

OpenStudy (anonymous):

if you split it up with seperate numerators and same denominators you get areally simply answer

OpenStudy (lgbasallote):

split it up \[\large \int \frac{x^2}{x} + \int \frac{3x}{x} - \int \frac{4}{x}\]

OpenStudy (anonymous):

Thanks guys